A section of boardwalk is made using 15 boards. Each board is 9 1/4 inches wide. The total width of the section is 144 inches. The spacing between each board is equal. What is the spacing between each board?
Found 2 solutions by fractalier, Cromlix:
Answer by fractalier(6550)

15 times 9.25 = 138.75 inches.
That leaves
144-138.75 = 5.25 inches for spaces.
Between the 15 boards are just 14 spaces, however.
So divide 5.25 by 14 and get
3/8" for each space.

Answer by Cromlix(4381)

Hi there,
14 boards each 9 1/4 inches wide.
14 x 9 1/4 = 138 3/4 inches
144 inches - 138 3/4 inches
= 5 1/4 inches
Spaces between 14 boards
= 13 spaces.
5 1/4 / 13 = 2/5 of an inch.
Hope this helps. :-)